# Making Images Real Again: A Comprehensive Survey on Deep Image Composition

### v5

@@ V-B Non-rendering based Methods
Other methods [203, 57, 92, 52] utilized paired training data (paired images with and without foreground shadow) to generate better shadow images. ShadowGAN [203] employed standard conditional GAN with reconstruction loss, local adversarial loss, and global adversarial loss to generate shadow for the inserted 3D foreground objects. Inoue et al. [57] developed a multi-task framework with two decoders accounting for depth map prediction and ambient occlusion map prediction respectively. ARShadowGAN [92] proposed an attention-guided residual network. The network predicts two attention maps for background shadow and occluder respectively, which are concatenated with composite image and foreground object mask to produce a residual shadow image. SGRNet [52] designed a two-stage shadow generation network. In the first stage, foreground features and background features are interacted using cross-attention to predict a shadow mask. In the second stage, they predict shadow parameters which are used to darken the input composite image. Then, the darkened image is combined with the input composite image with shadow matte. Meng et al. [107] adopted a similar two-stage pipeline and proposed to generate the shadow region by fusing multiple underexposure images. DMASNet [151] decomposed shadow mask prediction into box prediction and shape prediction, followed by attending relevant background shadow pixels to fill in the predicted shadow region. SGDiffusion [96] is the first work on shadow generation using diffusion model, which is built upon ControlNet [200] with extra intensity module to refine the shadow intensity.
@@ VI Generative Composition
As the diffusion models [128] pretrained on large-scale dataset [131] become popular in various image generation and editing tasks, generative image composition has attracted growing research interest. In contrast with previous methods which perform one or multiple sub-tasks sequentially, generative image composition is a combinatorial task which performs multiple sub-tasks (e.g., image blending, image harmonization, shadow generation) parallelly through one unified model. Given a foreground, a background, and a bounding box indicating the foreground placement, generative image composition targets at directly producing a realistic composite image with the foreground naturally and harmoniously merged into the background.
Generative image composition has certain overlap with object-guided image inpainting [183] and image customization [16, 187]. Object-guided image inpainting needs a mask to indicate the inpainted region, where the mask shape usually implies the target shape of inserted object. When the inpainted region is a bounding box, object-guided image inpainting is closer to generative image composition. However, in the strict sense, generative image composition expects to preserve the non-foreground pixels in the bounding box, which is different from object-guided image inpainting. Image customization is a very broad concept, which includes changing attributes and adding background for a specific object. Generative image composition can be deemed as a special case of image customization.
@@ VI-A Deep Learning Methods
The existing generative image composition methods can be divided into two groups: token-to-object methods and object-to-object methods. The first group of methods [79, 100] learn token-to-object mapping conditioned on the background information. Inspired by [129], they adapt a pretrained diffusion model to a subject by finetuning on a few reference images of this subject, so that a rare token is associated with this subject. For example, DreamEdit [79] proposed to finetune the diffusion model for one specific subject, and then perform iterative inpainting to combine foreground and background. DreamCom [100] proposed a more compact pipeline by finetuning text-guided image inpainting model. This group of methods can leverage multiple reference images of a single subject, but they are very time-consuming due to finetuning the pretrained diffusion model for each subject.
The second group of methods [183, 141, 205, 16, 191, 187] learn object-to-object mapping conditioned on the background information. They train a diffusion model on abundant pairs of foregrounds and backgrounds, so that it can be directly applied to a new pair of foreground and background at test time. In [183, 141], they construct massive training triplets of foregrounds, backgrounds, and ground-truth real images based on large-scale image datasets [69], in which the foregrounds are cropped from real images followed by color and geometry perturbation. Then, they adapt conditional diffusion model to this task. In particular, the background image, bounding box mask, and noisy image are concatenated as input, while the foreground is injected into the network via cross-attention. Kulal et al. [68] adopted a similar approach, but focused on human generation. Chen et al. [16] proposed to promote the foreground fidelity by using high-frequency information. The works [68, 16] pointed out that multi-view datasets and video datasets can help simulate more diverse and realistic geometry perturbation. Zhang et al. [191] proposed global-and-local fusion, in which shallow foreground features are used to enhance the details. Different works [16, 191, 187] also attempted to control image composition from different perspectives. For example, [16] provides the target shape of foreground object. [187] provides the target camera viewpoint of foreground object. [191] can selectively adjust the illumination and pose of foreground object to match the background.
Generative image composition methods perform multiple sub-tasks parallelly through a unified model, which is more concise than performing multiple sub-tasks sequentially. Nevertheless, the controllability is compromised in the meanwhile and undesirable changes could be made towards the foreground details.
@@ VI-B Datasets and Evaluation Metrics
Training diffusion model requires massive training triplets of foregrounds, backgrounds, and ground-truth real images. Previous works [183, 141] proposed to crop the foregrounds from real images and perturb the foregrounds (e.g., color transfer, geometric transformation), so that we can have perturbed foreground, masked background, and ground-truth real image. Chen et al. [16] argued that such simple perturbation cannot simulate real-world viewpoint change, so they resorted to multi-view image datasets and video datasets to get more reasonably perturbed foregrounds.
In real-world application scenarios, there exist no ground-truth images for a pair of foreground and background, so we cannot calculate the distance between generated image and ground-truth image. Therefore, previous works [183, 141, 16, 205] used FID [50] to measure the discrepancy between generated images and real images, quality score [42] to evaluate the authenticity of each image, and CLIP score [124] to measure the similarity between generated foreground and reference foreground.

### v6

@@ V-B Non-rendering based Methods
Other methods [225, 62, 101, 56] utilized paired training data (paired images with and without foreground shadow) to generate better shadow images. ShadowGAN [225] employed standard conditional GAN with reconstruction loss, local adversarial loss, and global adversarial loss to generate shadow for the inserted 3D foreground objects. Inoue et al. [62] developed a multi-task framework with two decoders accounting for depth map prediction and ambient occlusion map prediction respectively. ARShadowGAN [101] proposed an attention-guided residual network. The network predicts two attention maps for background shadow and occluder respectively, which are concatenated with composite image and foreground object mask to produce a residual shadow image. SGRNet [56] designed a two-stage shadow generation network. In the first stage, foreground features and background features are interacted using cross-attention to predict a shadow mask. In the second stage, they predict shadow parameters which are used to darken the input composite image. Then, the darkened image is combined with the input composite image with shadow matte. Meng et al. [116] adopted a similar two-stage pipeline and proposed to generate the shadow region by fusing multiple underexposure images. DMASNet [164] decomposed shadow mask prediction into box prediction and shape prediction, followed by attending relevant background shadow pixels to fill in the predicted shadow region. SGDiffusion [105] is the first work on shadow generation using diffusion model, which is built upon ControlNet [222] with extra intensity module to refine the shadow intensity. [189, 165] also train conditional diffusion model for shadow generation. [231, 208] first predict coarse shadow mask and then feed the shadow mask to diffusion model.
@@ V-C Datasets and Evaluation Metrics
Instead of constructing synthetic datasets [56, 105] by removing the shadows, [189] constructed real-world dataset by taking photos with object (factual image) or without object (counterfactual image). However, this approach to construct dataset is very costly and labor-intensive.
@@ VI Generative Composition
As the diffusion models [140] pretrained on large-scale dataset [142] become popular in various image generation and editing tasks, generative image composition (object compositing) has attracted growing research interest. In contrast with previous methods which perform one or multiple sub-tasks sequentially, generative image composition is a combinatorial task which performs multiple sub-tasks (e.g., image blending, image harmonization, shadow generation) parallelly through one unified model. Given a foreground, a background, and a bounding box indicating the foreground placement, generative image composition targets at directly producing a realistic composite image with the foreground naturally and harmoniously merged into the background.
Generative image composition has certain overlap with object-guided image inpainting [202] and image customization [209]. Their differences are claimed as follows. 1) Object-guided image inpainting needs a mask to indicate the inpainted region, where the mask shape usually implies the target shape of inserted object. When the inpainted region is a bounding box free of shape information, object-guided image inpainting is closer to generative image composition. However, strictly speaking, generative image composition expects to preserve the non-foreground pixels in the bounding box, which is different from object-guided image inpainting. Moreover, generative image composition aims to generate shadow and reflection for the foreground object without box or shape constraint, which is also different from object-guided image inpainting. 2) Image customization is a very broad concept, which includes changing attributes and adding background for a specific object. Generative image composition can be deemed as a special case of image customization.
@@ VI-A Deep Learning Methods
The existing generative image composition methods can be divided into two groups: training-free methods and training-based methods.
The first group of methods [50, 110, 183] utilizes off-the-shelf foundation generation model, which does not require training or finetuning. They aim to generate high-quality composite images by manipulating the foreground and background elements (e.g., feature, attention) through the denoising process.
The second group of methods [202, 152, 227, 213, 209, 19, 88, 109, 163, 228, 189] learn object-to-object mapping conditioned on the background information. They train a diffusion model on abundant pairs of foregrounds and backgrounds, so that it can be directly applied to a new pair of foreground and background at test time. In the testing period, if a few images containing the foreground object are available, the pretrained model can also be finetuned on these images for better performance.
In the pioneering works [202, 152], they construct massive training triplets of foregrounds, backgrounds, and ground-truth real images based on large-scale image datasets [75], in which the foregrounds are cropped from real images followed by color and geometry perturbation. Then, they adapt conditional diffusion model to this task. In particular, the background image, bounding box mask, and noisy image are concatenated as input, while the foreground is injected into the network via cross-attention. Kulal et al. [74] adopted a similar approach, but focused on human generation. Some subsequent methods focuses on enhancing the ability of detail preservation. For example, Zhang et al. [213] proposed global-and-local fusion, in which shallow foreground features are used to enhance the details. Chen et al. [19] extracted high-frequency information for better detail preservation. Different works [213, 209] also attempted to control image composition from different perspectives. For example, [209] provided the target camera viewpoint of foreground object. [213] can selectively adjust the illumination and pose of foreground object to match the background. Some methods [53, 86, 230] aimed to insert the object into any reasonable place in the background image without the provided bounding box. Some methods [230, 190] explored generating plausible shadow and reflection for the inserted foreground without the spatial constraint of bounding box.
@@ VI-B Datasets and Evaluation Metrics
Training diffusion model requires massive training triplets of foregrounds, backgrounds, and ground-truth real images. Previous works [202, 152] proposed to crop the foregrounds from real images and perturb the foregrounds (e.g., color transfer, geometric transformation), so that we can have perturbed foreground, masked background, and ground-truth real image. The multi-view datasets and video datasets can also be used to simulate more diverse and realistic geometry perturbation.
In real-world application scenarios, there exist no ground-truth images for a pair of foreground and background, so we cannot calculate the distance between generated image and ground-truth image. Therefore, previous works [202, 152, 227] used FID [54] to measure the discrepancy between generated images and real images, quality score [45] to evaluate the authenticity of each image, and CLIP score [135] to measure the similarity between generated foreground and reference foreground.